OpenStudy (anonymous):

Suppose y varies directly with x, and y = 14 when x = -4. What is the value of y when x = -6? Can anyone help me with this? Like explain it because I don't understand it like how does this work?

OpenStudy (phi):

y varies directly with x means y = k*x where k is some number your first step is find k start with y = 14 when x = -4 replace y with 14 and x with -4 in the y=k*x equation: 14= k* -4 can you "solve for k"? (divide both sides by -4)

OpenStudy (anonymous):

14/-4 = k*-4/-4 -3.5 = k? right?

OpenStudy (anonymous):

Wait but the choices on it says a. -21 b. 10 c. 18 d. 21

OpenStudy (phi):

so far, so good. k= -3.5 now you know k, so "direct variation" means y = -3.5 x to find y, replace x with a number. What is the value of y when x = -6?

OpenStudy (phi):

The idea is multiply x times -3.5. If x is -6, that means multiply -6*-3.5 to get the y

OpenStudy (anonymous):

okay so it would be 21?

OpenStudy (anonymous):

so d?

OpenStudy (phi):

yes. To summarize direct variation means y = k x find k (you need an x and a y value to do that) once you know k, you can find y for any other x they give you.

OpenStudy (anonymous):

Thanks you helped me a lot! :)